OpenStudy (anonymous):

solve the quadratic equation 12x^2-3x-9=0 by factoring

OpenStudy (mertsj):

1. factor out 3

OpenStudy (anonymous):

These questions should be done at wolframalpha to check your answers http://www.wolframalpha.com/input/?i=12x^2-3x-9%3D0

OpenStudy (whpalmer4):

No, better that you check your answer by multiplication; most people doing factoring problems can use the practice, in my experience. Watching the results of the multiplication come together also give you a better eye for factoring.

OpenStudy (anonymous):

12x^2-3x+9=0 >4x^2-x+3=0 >4x^2-4x+3x+3=0 >4x(x-1)+3(x-1)=0 >(x-1)(4x-3)=0 therefore,x-1=0 or,4x-3=0>x=3/4 >x=1

OpenStudy (anonymous):

12x^2-3x-9=0 (3)(4x^2-x-3)=0 (3)(x-1)(4x+3)=0 (3x-1)(4x+3)=0 Therefore; x=1/3,-3/4

OpenStudy (whpalmer4):

@Ezack \[3(x-1)(4x+3) \ne (3x-1)(4x+3)\]Also, no need to incorporate the 3 when solving. \[a(x-1)(4x+3)\]has the same roots for any value of \(a\) that is not 0.

OpenStudy (anonymous):

@whpalmer4, thanks for the correction, it's (3x-3) which means x=1, -3/4

OpenStudy (whpalmer4):

@raselazam \[12x^2-3x-9=0\]\[4x^2-x-3=0\]\[4*-3=-12\]-4,3 are factors of -12 that add to -1 \[(4x^2-4x) +(3x-3) = 0\]\[4x(x-1)+3(x-1) = 0\]\[(4x+3)(x-1) = 0\]\[4x+3=0\]\[4x=-3\]\[x=-\frac{3}{4}\]\[x-1=0\]\[x=1\]So solutions are \[x=1,-\frac{3}{4}\]

OpenStudy (anonymous):

you are right actually the answer is 1 and -3/4.but unconsciously I give negative sign ,so the answer 3/4 is wrong.But how can you tell that the process is wrong? how? would you please explain?

OpenStudy (whpalmer4):

Reading your work, you changed some signs... >4x^2-4x+3x+3=0 >4x(x-1)+3(x-1)=0 >(x-1)(4x-3)=0 in the first line, you've got 3x+3 but it should be 3x-3 (harmless error as you fix it in the next line but then going from 2nd to 3rd line, you change (4x+3) to (4x-3) which isn't harmless...

OpenStudy (whpalmer4):

When I factor anything, I almost always multiply the factors out to make sure I end up with the polynomial I started with... \[(x-1)(4x-3) = 4x^2 - 3x -4x -3 = 4x^2 -7x -3\]which isn't what I started with...

OpenStudy (anonymous):

Thank you.I found out my mistake.But in this case, is there any way to solve this question applying factoring method?

OpenStudy (whpalmer4):

I thought we did solve this problem by factoring...I don't understand your question.

OpenStudy (anonymous):

I just want to know from you is there any alternative way of factorising that is helpful to find out the roots of the equation.